# Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable success across diverse tasks, but, due to single-token generation at each decoding step, their inference processes are hindered by substantial time and energy demands. While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution. Although this enhances the speed, it does not improve the output quality. In contrast, our work simultaneously boosts inference speed and improves the output effectiveness. We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and raising task performance. However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens. Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD. MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also increases the decoding efficiency over conventional speculative decoding. Theoretically, we demonstrate that MTAD closely approximates exact MTJD with a bounded error. Empirical evaluations across various tasks reveal that MTAD improves downstream performance by 25% compared to standard single-token sampling. Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than vanilla speculative decoding methods. These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more productive and high-performance deployment of LLMs. [^1] We release our code at https://github.com/ZongyueQin/MTAD
@@ 1 Introduction
    We conducted comprehensive evaluations across various tasks, demonstrating that MTAD improves downstream performance by 25% compared to standard single-token sampling, while, at the same time, realizing a 1.42$\times$ speed-up and paring energy consumption by 1.54$\times$ compared to vanilla speculative decoding methods.
@@ 3.2 Multi-Token Assisted Decoding
Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to produce multiple draft sequences and by applying tree-based attention (Miao et al., 2023) for simultaneous verification. However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant. Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above. Therefore, we leave a more detailed exploration of this extension as future work.
@@ 5.2 Performance of Multi-Token Assisted Decoding
We first observe that MTAD is the most efficient among all baselines in terms of both energy and time. On average, MTAD is 17.2% faster than the most efficient lossless baseline, MCSS, while consuming 6.1% less energy. Compared to lossy decoding algorithms, it attains 27.2% (15.1%) higher speed and 17.9% (14.1%) lower energy consumption than BiLD (typical decoding), respectively. Interestingly, despite utilizing only a single draft sequence, MTAD outperforms baselines that employ multiple draft sequences, such as Spectr, MCSS, and SpecInfer. In these methods, verification terminates immediately upon rejecting a token. In contrast, MTAD continues verification even after a rejection, searching for future tokens that may still pass. This mechanism results in a greater acceptance length per iteration than the baselines.
@@
Next, we compare the downstream performance of different decoding algorithms. Notice that while lossless decoding algorithms theoretically sample from the target distribution, they exhibit slight variations in downstream performance. This discrepancy arises because, despite preserving the original distribution, differences in the sampling process prevent them from generating identical sequences even when the random seed is fixed. Furthermore, we observe that lossy decoding algorithms can reach higher downstream performance at the expense of efficiency. This suggests that all lossy decoding methods can trade off efficiency for performance by adjusting verification strictness. Most notably, MTAD consistently achieves the highest downstream performance. On average, it surpasses lossless decoding algorithms by 25.0%, BiLD by 11.8%, and typical decoding by 9.9%. These results confirm our claim that MTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
@@
BiLD Typical SpD Spectr SpecInfer MCSS MTAD
@@
tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 24.8
J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 7.6
pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 38.4
@@
tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 25.6
J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 7.6
pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 22.0
@@
tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 23.5
J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 9.2
Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0
@@
tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 $\downarrow$ 33.3
J/token 8.2 7.9 7.5 7.1 8.1 8.0 7.8
Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 35.0
@@
tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 $\downarrow$ 29.8
J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2
score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40
@@
tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2
J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0
score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75
@@ C.1 Additional Datasets and Model Family
Table 10: Number of Generated Tokens Per Iteration Across Thresholds
Threshold
0.1 0.3 0.5 0.7 0.9
Llama-3 4.72 4.43 4.12 3.93 3.64
Llama-3-Instruct 4.74 4.51 4.28 4.06 3.79
C.2 Ablation Studies on Acceptance Threshold
Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes stricter, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 3, the efficiency indeed drops when $\tau$ increases. However, the perplexity rises when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence MTAD’s perplexity is similar to the perplexity of multinomial sampling. Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
In addition, Table 10 exhibits how the number of accepted tokens change with thresholds. We can see that the number of accepted tokens grows smaller when the threshold increases, which explains the speed degradation.
C.3 Ablation Studies on Number of Beams
In Table 11, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance. Increasing the number of beams boosts the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency. But we can see that the increment slows down when the number of beams is too large. In addition, when the number of beams is too large, the inference cost of the small model will become too high.
Table 11: Effect of beam count on inference performance (ChatGPT-Prompts dataset). Arrows indicate optimal direction ($\uparrow$=higher better, $\downarrow$=lower better).
Number of Beams
2 4 6 8
Llama-2 Speed (tok/s) $\uparrow$ 55.9 59.9 60.2 61.3
Energy (J/tok) $\downarrow$ 2.43 2.25 2.22 2.20
Perplexity $\downarrow$ 2.44 2.12 2.14 2.10
OPT Speed (tok/s) $\uparrow$ 51.0 54.1 54.3 55.9
Energy (J/tok) $\downarrow$ 2.50 2.32 2.36 2.30
Perplexity $\downarrow$ 3.63 3.16 3.42 3.19
C.4 Ablation study of top-k and top-p sampling
@@
C.5 Resutlts with OPT-30B and Llama-2-70B
@@
C.6 Additional Experiments on CNNDM and Spider
Table 14 and Table 15 reveal the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset. We can see that MTAD is consistently more effective and has a faster decoding speed.
Table 14: Comparison of ROUGE-L Scores and Tokens per Second under Different Fine-Tuning Conditions on CNNDM
No Fine-Tune Fine-Tune 68M Fine-Tune Both
Method ROUGE-L Tokens/sec ROUGE-L Tokens/sec ROUGE-L Tokens/sec
SpeD 0.114 37.7 0.114 20.4 0.164 24.3
MTAD 0.118 44.2 0.121 25.0 0.168 27.1
Table 15: Comparison of Execution Accuracy (EA) and Tokens per Second under Different Fine-Tuning Conditions on Spider
No Fine-Tune Fine-Tune 68M Fine-Tune Both
Method EA Tokens/sec EA Tokens/sec EA Tokens/sec
SpD 11.5 28.5 11.5 27.8 16.3 25.6
MTAD 13.0 30.3 14.8 32.3 18.3 29.4
C.7 Visualization of perplexity and output quality
@@
C.8 Correlation between Energy and Speed
Figure 5 shows there is a correlation between speed and energy, whether we consider the entire table or focus on a specific dataset and model. For fairness, all methods for a given dataset and model were run on the same machine nodes. However, for a fixed method (e.g., Spectr), experiments on different datasets and models might be conducted on different nodes (all equipped with L40 GPUs). We did notice that the same configuration run on different machines may have varied energy consumption. This variation introduces some randomness, which could make the correlation appear less consistent across datasets and models.
(a) group by dataset and model
(b) group by methods
Figure 5: Correlation between speed and energy

### v4

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@ 3.2 Multi-Token Assisted Decoding
3.3 Multi Candidate Verification
As illustrated in Figure 3, the intermediate results of beam sampling naturally form a tree structure, where each layer contains $b$ nodes corresponding to the $b$ intermediate beams generated at that step. In vanilla MTAD, only the final output sequence is retained and verified by the target model, while all other intermediate beams are discarded. However, these discarded beams may in fact have higher likelihood under the target model. To address this limitation, we propose an enhanced version of MTAD that leverages all intermediate beams during verification. This optimization introduces two key benefits: (1) Improved efficiency. By incorporating more candidates into the verification process, the probability of accepting tokens at each step increases, which leads to faster decoding. (2) Enhanced output quality. Among the accepted candidates, we can select the one with the highest target likelihood, potentially yielding better generations than vanilla MTAD.
To efficiently verify all intermediate beams, we adopt the tree attention mechanism introduced by Miao et al. (2023), which allows the target model to compute the conditional likelihood of every token in the draft tree in a single forward pass. Specifically, at each layer $i$, we evaluate the target likelihood $p(x_{t+1:t+i}^{(j)}\mid x_{1:t})$ for each beam $j\in{1,\dots,b}$.
In vanilla MTAD, a beam sequence $x_{t+1:t+i}$ is accepted if its likelihood ratio $\frac{p(x_{t+1:t+i}\mid x_{1:t})}{q(x_{t+1:t+i}\mid x_{1:t})}\geq\tau$, which is a reasonable criterion when $x_{1:t}$ corresponds to a prefix of the final output. However, when verifying all intermediate beams—particularly with a large beam width $b$, many beams may have low draft likelihood $q(x_{t+1:t+i}^{(j)}\mid x_{1:t})$, making the denominator small and potentially resulting in the acceptance of low-quality candidates.
To mitigate this issue, we revise the acceptance criterion: instead of comparing each beam against its own draft likelihood, we normalize all beams against the draft likelihood of the highest-likelihood final beam. Specifically, we accept a candidate $x_{t+1:t+i}^{(j)}$ if
$\frac{p(x_{t+1:t+i}^{(j)}|x_{1:t})}{q(x_{t+1:t+i}^{**}|x_{1:t})}\geq\tau$ (9)
where $x_{t+1:t+\gamma}^{**}$ is the output sequence of the beam sampling, which is the beam with the highest draft likelihood at the last step.
Once all beams in the tree are verified, we select the longest accepted sequence and generate one additional token from the target model $p$ to be the output tokens at this iteration. If there are multiple accepted sequences with same lengths, the sequence with the highest target likelihood is chosen.
This variant, which we call multi-candidate MTAD (MMTAD), improves both the robustness and effectiveness of decoding. The following theorem demonstrates that MMTAD achieves a higher acceptance rate at each decoding step compared to vanilla MTAD, which directly translates to better decoding efficiency.
Theorem 3.7.
MMTAD has a higher expected accepted sequence length than vanilla MTAD.
Proof.
Let $x^{**}_{t+1:t+\gamma}$ denote the output sequence of beam sampling. The accepted length of vanilla MTAD is $\max\{i:\frac{p(x^{**}_{t+1:t+i}|x_{1:t})}{q(x^{**}_{t+1:t+i}|x_{1:t})}\}\geq\tau\}$. Notice that MMTAD also verifies $x^{**}$ and the acceptance criterion does not change for any $x^{**}_{t+1:t+i}$. Meanwhile, MMTAD verifies more candidates which might pass the verification when $x^{**}_{t+1:t+i}$ fails. Therefore, MMTAD has a higher expected accepted length than MTAD. ∎
Additionally, the next theorem confirms that MMTAD also has bounded error on the perplexity ratio.
Theorem 3.8.
Let $x_{1:\Gamma_{N}}^{\text{multi}}$ be the sequence generated by multi-candidate MTAD (MMTAD), and $x_{1:\Gamma_{N}}^{*}$ be the sequence generated by deterministic exact MTJD. Under the same assumption as Theorem 3.5, i.e., $\forall x_{1:t}\in\mathcal{X}$, $|\log p(x|x_{1:t})-\log q(x|x_{1:t})|_{\infty}\leq U$, the perplexity ratio of MMTAD is bounded as:
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}}^{\text{multi}})}{PPL_{p}(x_{1:\Gamma_{N}}^{*})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (10)
The proof is provided in the Appendix A.
Although Theorem 3.8 indicates that MMTAD has a slightly looser perplexity ratio bound than vanilla MTAD, due to the additional $(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}$ term, in practice, MMTAD often yields higher-quality outputs. This is because it selects the longest verified sequence at each step, and in the case of ties, chooses the one with the highest target model likelihood. This strategy prioritizes more promising candidates and typically results in better generations than vanilla MTAD.
(a) The draft tree of beam sampling
(b) tree attention mask
(c) MMTAD verification
Figure 3: Illustration of MMTAD: (a) All intermediate beams of beam sampling naturally form a tree. Vanilla MTAD only verify the output beam (yellow blocks), MMTAD verify all the beams. (b) MMTAD utilizes tree attention to efficiently compute the target likelihood of each beam. (c) MMTAD selects the longest accepted sequence with the highest target likelihood to return.
@@ 5.2 Performance of Multi-Token Assisted Decoding
We first observe that MMTAD is the most efficient among all baselines in terms of both energy and time. On average, MMTAD is 21.1% faster than the most efficient lossless baseline, MCSS, while consuming 23.4% less energy. Compared to lossy decoding algorithms, it attains 46.8% (31.5%) higher speed and 30.4% (28.3%) lower energy consumption than BiLD (typical decoding), respectively. Interestingly, despite utilizing only a single draft sequence, MTAD outperforms baselines that employ multiple draft sequences, such as Spectr, MCSS, and SpecInfer. In these methods, verification terminates immediately upon rejecting a token. In contrast, MTAD continues verification even after a rejection, searching for future tokens that may still pass. This mechanism results in a greater acceptance length per iteration than the baselines.
@@
Next, we compare the downstream performance of different decoding algorithms. Notice that while lossless decoding algorithms theoretically sample from the target distribution, they exhibit slight variations in downstream performance. This discrepancy arises because, despite preserving the original distribution, differences in the sampling process prevent them from generating identical sequences even when the random seed is fixed. Furthermore, we observe that lossy decoding algorithms can reach higher downstream performance at the expense of efficiency. This suggests that all lossy decoding methods can trade off efficiency for performance by adjusting verification strictness. Most notably, MMTAD consistently achieves the highest downstream performance. On average, it surpasses lossless decoding algorithms by 42.7%, BiLD by 27.3%, and typical decoding by 24.8%. These results confirm our claim that MMTAD offers superior effectiveness compared to conventional decoding methods that rely solely on single-token distributions.
@@
BiLD Typical SpD Spectr SpecInfer MCSS MTAD MMTAD
@@
tokens/s $\uparrow$ 17.4 21.7 22.2 23.8 22.8 23.7 28.2 29.7
J/token $\downarrow$ 10.0 8.1 7.8 7.8 7.9 7.8 5.6 5.5
pass@1 $\uparrow$ 37.8 35.9 32.9 32.9 31.0 32.0 43.2 45.1
@@
tokens/s $\uparrow$ 19.6 22.5 22.2 24.4 22.5 23.8 27.5 29.1
J/token $\downarrow$ 9.7 8.9 8.9 8.9 8.1 7.9 6.1 6.1
pass@1 $\uparrow$ 19.5 20.0 15.9 16.0 17.7 17.0 26.8 28.0
@@
tokens/s $\uparrow$ 20.1 22.3 19.6 22.4 21.1 21.7 22.8 25.5
J/token $\downarrow$ 10.2 9.5 10.5 9.6 10.2 10.0 8.1 7.7
Acc $\uparrow$ 35.0 42.0 36.0 35.5 37.0 35.0 44.0 54.0
@@
tokens/s $\uparrow$ 23.3 32.3 31.1 32.1 32.6 32.7 37.9 43.4
J/$\downarrow$ 8.2 7.9 7.5 7.1 8.1 8.0 6.0 5.5
Acc $\uparrow$ 30.5 29.5 21.5 23.0 21.5 24.0 39.0 40.0
@@
tokens/s $\uparrow$ 25.9 23.4 26.0 26.2 26.3 26.8 29.8 32.9
J/token $\downarrow$ 10.8 12.2 10.0 9.9 10.0 9.9 9.2 7.5
score $\uparrow$ 4.15 4.26 4.10 4.11 4.01 4.02 4.40 4.39
@@
tokens/s $\uparrow$ 24.5 22.3 24.1 24.5 24.5 25.7 28.2 29.8
J/token $\downarrow$ 11.5 12.4 11.0 11.6 11.7 11.1 10.0 10.1
score $\uparrow$ 3.41 3.24 3.39 3.41 3.35 3.36 3.75 3.75
5.3 Ablation Studies
5.3.1 Draft Sequence Length
We investigate the impact of the draft sequence length $\gamma$ on the performance of MMTAD. Figure 4 presents results for decoding speed, block efficiency (i.e., the average number of tokens generated per iteration), and output perplexity using the Llama-3-8B-Instruct model on the Spider dataset, with $\gamma\in\{3,4,5,6,7,8,9,10\}$. As $\gamma$ increases, block efficiency consistently improves. However, decoding speed saturates once $\gamma$ reaches 7. This is due to the growing computational overhead associated with generating and verifying longer draft sequences, which offsets the gains from improved block efficiency. Meanwhile, as shown in Figure 4(c), output perplexity remains stable across different values of $\gamma$.
(a) $\gamma$ vs. Speed
(b) $\gamma$ vs. Block efficiency
(c) $\gamma$ vs. Perplexity
Figure 4: Performance of MMTAD when draft length $\gamma\in\{3,4,5,6,7,8,9,10\}$.
5.3.2 Beam Sampling Width
Next, we study the effect of beam width $b$ used by the draft model when generating draft sequences for MMTAD. As shown in Figure 5, increasing $b$ leads to a slight improvement in block efficiency. This is because having more candidate beams increases the likelihood that more tokens will be accepted during verification. Additionally, output perplexity shows a slight decrease as $b$ increases, since MMTAD selects the longest accepted sequence with the highest likelihood under the target model. We also observe that decoding speed initially improves with larger $b$, owing to gains in block efficiency, but plateaus once $b$ reaches 5. This is due to the growing cost of verifying a larger number of candidate beams, which offsets the speedup from increased acceptance.
(a) $b$ vs. Speed
(b) $b$ vs. Block efficiency
(c) $b$ vs. Perplexity
Figure 5: Performance of MMTAD when beam width $b\in\{2,3,4,5,6\}$.
5.3.3 Acceptance Threshold
Finally, we examine the impact of the acceptance threshold $\tau$. As shown in Figure 6, increasing $\tau$ imposes a stricter acceptance criterion, leading to lower block efficiency and reduced decoding speed. This is expected, as fewer draft sequences satisfy the higher acceptance threshold. On the other hand, output perplexity decreases as $\tau$ increases, since stricter acceptance favors higher-confidence predictions. However, this improvement saturates as $\tau$ approaches 1, with diminishing returns in perplexity reduction.
(a) $\tau$ vs. Speed
(b) $\tau$ vs. Block efficiency
(c) $\tau$ vs. Perplexity
Figure 6: Performance of MMTAD when acceptance threshold $\tau\in\{0.1,0.3,0.5,0.7,0.9\}$.
@@ Appendix A Proof
A.6 Proof of Theorem 3.8
Proof.
Let $x^{**}$ be the output sequence of beam sampling. We have
$\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\geq\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{\max_{x^{{}^{\prime}}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}}q(x^{{}^{\prime}}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}$ (34)
$\geq$ $\frac{(1-\epsilon)p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x^{**}_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\quad(\text{assumption of beam sampling error})$
$\geq$ $(1-\epsilon)\tau$
With the same procedure in the proof of Lemma A.1, we have
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\gamma_{N}})}{PPL_{q}(x_{1:\gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}$ (35)
Then, with the same procedure to prove Theorem 3.5, we have
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}(1-\epsilon)^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (36)
@@ C.1 Additional Datasets and Model Family
C.2 Ablation study of top-k and top-p sampling
@@
C.3 Resutlts with OPT-30B and Llama-2-70B
@@
C.4 Visualization of perplexity and output quality